# Transitivity of the pure Hurwitz classes of quadratic post-critically finite polynomials

### v1

@@ 3. Portrait functions
    Create disjoint and intersecting pre-periods
For each function, we specify by which element $\phi$ of $\operatorname{Mod}(S^{2},M_{f})$ we pre-compose $f$ by to modify the portrait as desired. We then prove that the proposed composition modifies the portrait as desired.
Function 1 (Increase a cycle length).
Take as input a marked branched cover $f$ with marked set $M$ and portrait $\Gamma$. Let $f^{\ell}(C_{j})$ be a point in the cycle of the critical point $C_{j}$. Let $\star$ denote a fixed point of $f$. Let $\phi$ be a half twist that swaps $\star$ with $f^{\ell}(C_{j})$. Output the marked branched cover $f\phi$, the marked set $M\cup\{\star\}$, the portrait of $f\phi$, and the mapping class $\phi$. The cycle of $C_{j}$ in the portrait of $f\phi$ has length one more than the cycle of $C_{j}$ in the portrait of $f$. The branched cover $f\phi$ only affects the length of the cycle, and no other portrait features of $\Gamma_{f}$. We refer to the change in the portrait that results from applying $\phi$ as “applying Function 1 to the point $f^{\ell}(C_{j})$.”
Proof.
The map $f\phi$ maps $f^{\ell}(C_{j})$ to $\star$, and $\star$ to $f^{\ell+1}(C_{j})$. Therefore the portrait for $f\phi$ has one more point than $\Gamma_{f}$. Because $\phi$ is supported on a neighborhood of an arc between $f^{\ell}(C_{j})$ and $\star$, no other cycles or pre-periods of $\Gamma_{f}$ change under the map $\phi$. That is: the length of the cycle increases by one and all other portrait features remain unchanged. ∎
Function 2 (Decrease a cycle length).
Take as input a marked branched cover $f$ with marked set $M$ and portrait $\Gamma$. If the length of the cycle of $C_{j}$ is greater than one, let $f^{\ell}(C_{j})$ and $f^{\ell+1}(C_{j})$ be (distinct) post-critical points in the cycle of $C_{j}$. Suppose that $f^{\ell+1}(C_{j})$ is not a critical point and $\Gamma$ satisfies at least one of the following conditions:
  1. (1)
    $f^{\ell}(C_{j})$ is not a critical point,
  2. (2)
    the component of $\Gamma$ containing $C_{j}$ is a cycle, or
  3. (3)
    $f^{\ell+1}(C_{j})\neq f^{r+k-1}(C_{j})$.
Then let $\phi$ be a half twist that swaps $f^{\ell}(C_{j})$ and $f^{\ell+1}(C_{j})$. Output the marked branched cover $f\phi$, the marked set $M$, the portrait of $f\phi$, and the mapping class $\phi$. The cycle of $C_{j}$ in the portrait of $f\phi$ has length one less than the the length of the cycle of $C_{j}$ in the portrait of $f$.
The branched cover $f\phi$ only affects the length of the cycle, and no other portrait features of $\Gamma_{f}$. We refer to the change in the portrait that results from applying $\phi$ as “applying Function 2 to the point $f^{\ell}(C_{j})$.”
Proof.
Function 3 (Increase a pre-period length).
Take as input a marked branched cover $f$ with marked set $M$ and portrait $\Gamma$ such that at least one critical point is strictly pre-periodic. Let $\star$ denote a fixed point of $f$. We can either increase the length of the pre-period of $C_{j}$ that is disjoint from the pre-period of $C_{j^{\prime}}$ with $j^{\prime}\neq j$ or we can increase the length of the intersection of the pre-periods of $C_{1}$ and $C_{2}$.
  1. (1)
    Let $\phi_{1}$ be a half twist that swaps $C_{j}$ and $\star$. Output the marked branched cover $f\phi$, the marked set $M\cup\{\star\}$, the portrait of $f\phi$, and the mapping class $\phi$. The pre-period of $C_{j}$ in the portrait of $f\phi$ has one more point than the pre-period of $C_{j}$ in the portrait of $f$. In particular, the additional point occurs in the portion of the pre-period of $C_{j}$ in the portrait of $f\phi$ that is disjoint from the pre-period of the other critical point. We refer to the change in the portrait that results from applying $\phi$ as “applying Function 3(1) to the point $C_{j}$.”
  2. (2)
    Suppose the pre-periods of $C_{1}$ and $C_{2}$ intersect (non-trivially). Let $f^{\ell}(C_{1})$ be in the intersection of the pre-periods of $C_{1}$ and $C_{2}$. Let $\phi$ be a half twist that swaps $f^{\ell}(C_{1})$ and $\star$. Output the marked branched cover $f\phi$, the marked set $M\cup\{\star\}$, the portrait of $f\phi$, and the mapping class $\phi$. Then the intersection of the pre-periods of $C_{1}$ and $C_{2}$ in the portrait of $f\phi$ has one more point than the intersection of the pre-periods of $C_{1}$ and $C_{2}$ in the portrait of $f$. We refer to the change in the portrait that results from applying $\phi$ as “applying Function 3(2) to the point $f^{\ell}(C_{i})$.”
The branched cover $f\phi$ only affects the length of the pre-period, and no other portrait features of $\Gamma_{f}$.
Proof.
Let $\phi$ denote the twist which swaps $f^{\ell}(C_{j})$ with $\star$ (where $\ell$ may equal 0 in Case (1)). Applying $f$ after the twist $\phi$ results in a map where $f^{\ell}(C_{j})$ maps to $\star$ and $\star$ maps to $f^{\ell+1}(C_{j})$. Because $\phi$ is supported on a neighborhood of an arc between $f^{\ell}(C_{j})$ and $\star$, no other portrait features of $\Gamma_{f}$ change under the map $f\phi$. ∎
Function 4 (Decrease a pre-period length).
Take as input a marked branched cover $f$ with marked set $M$ and portrait $\Gamma$ that has at least one pre-periodic critical point $C_{j}$ with pre-period length at least 3. We can either decrease the length of the pre-period of $C_{j}$ that is disjoint from the pre-period of $C_{j^{\prime}}$ with $j^{\prime}\neq j$ or we can decrease the length of the intersection of the pre-periods of $C_{1}$ and $C_{2}$.
  1. (1)
    Suppose the length of the pre-period of $C_{j}$ under $f$ that is disjoint from the pre-period of the other critical point under $f$ has length at least 3. Let $f^{\ell}(C_{j})$ and $f^{\ell+1}(C_{j})$ be (distinct) post-critical points in the pre-period of $C_{j}$ that is disjoint from the pre-period of the other critical point (if it exists). Let $\phi$ be a half twist that swaps $f^{\ell}(C_{j})$ and $f^{\ell+1}(C_{j})$. Output the marked branched cover $f\phi$, the marked set $M\cup\{\star\}$, the portrait of $f\phi$, and the mapping class $\phi$. Then the length of the pre-period of $C_{j}$ under $f\phi$ that is disjoint from the pre-period of the other critical point under $f\phi$ is one less than the length of the pre-period of $C_{j}$ under $f$ that is disjoint from the pre-period of the other critical point under $f$.
  2. (2)
    Suppose the length of the intersection of the pre-periods of $C_{1}$ and $C_{2}$ has length at least 3. Let $f^{\ell}(C_{j})$ and $f^{\ell+1}(C_{j})$ be (distinct) post-critical points in the intersection of the pre-periods of $C_{1}$ and $C_{2}$ such that neither $f^{\ell}(C_{j})$ nor $f^{\ell+1}(C_{j})$ is a critical point. Let $\phi$ be a half twist that swaps $f^{\ell}(C_{j})$ and $f^{\ell+1}(C_{j})$. Output the marked branched cover $f\phi$, the marked set $M\cup\{\star\}$, the portrait of $f\phi$, and the mapping class $\phi$. The length of the intersection of the pre-periods of $C_{1}$ and $C_{2}$ under $f\phi$ has one fewer point than the intersection of the pre-periods of $C_{1}$ and $C_{2}$ under $f$.
In both cases, the set $M\cup\{\star\}$ will contain points that are not in the portrait of $f\phi$. The branched cover $f\phi$ only affects the length of the pre-period, and no other portrait features of $\Gamma_{f}$. We refer to the change in the portrait that results from applying $\phi$ as “applying Function 4 to the point $f^{\ell}(C_{j})$.”
Proof.
The branched cover $f\phi$ maps $f^{\ell}(C_{j})$ to $f^{\ell+2}(C_{j})$ and maps $f^{\ell+1}(C_{j})$ to itself. Because $\phi$ can be chosen to be supported on a neighborhood of an arc connecting $f^{\ell}(C_{j})$ and $f^{\ell+2}(C_{j})$ that is disjoint from all other critical and post-critical points of $f\phi$, no other portrait features of $\Gamma_{f}$ change under $f\phi$. ∎
Function 5 (Move $C_{2}$ within the cycle of $C_{1}$).
Take as input a marked branched cover $f$ with marked set $M$ and associated portrait $\Gamma$ with one component such that one critical point is strictly pre-periodic and the other critical point is periodic. Keeping to our conventions, assume that $C_{2}=f^{q}(C_{1})$. Let $r$ be the length of the pre-period of $C_{1}$ so that $f^{r}(C_{1})$ is the first periodic point in the orbit of $C_{1}$. Because $C_{2}$ is a critical point, $f(C_{2})$ is not equal to $f^{r}(C_{1})$. Let $q-r$ be the distance between $f^{r}(C_{1})$ and $C_{2}$ and $k+r-q$ be the distance between $C_{2}$ and $f^{r}(C_{1})$. We keep the cycle length $k$ constant, but increase $q$ or decrease $q$.
  • (E)
    Decrease $q$: Suppose that $C_{2}$ is not the first vertex in the cycle (in which case $q$ cannot be decreased). Let $x=f^{r+1}(C_{1})$ be the second post-critical point in the cycle (which could be $C_{2}$). Let $\diamond$ be the element of $f^{-1}(x)$ that is not post-critical. Let $\phi$ be the half twist that swaps $\diamond$ with $f^{r-1}(C_{1})$. Output $f\phi$, the marked set $M\cup\{\diamond\}$, the portrait of $f\phi$ (in this case, not all points in $M$ are vertices in the portrait of $f\phi$), and the mapping class $\phi$. Then $C_{2}$ occurs one position earlier in the cycle of $C_{1}$ under $f\phi$ than $C_{2}$ occurs in the cycle of $C_{1}$ under $f$, so $q-r$ decreased by one and the cycle length remained constant.
  • (L)
    Increase $q$: Suppose that $C_{2}$ is not the second-last (or last) point in the cycle (it cannot be last, so $q$ cannot be increased if $C_{2}$ is the penultimate point of the cycle). Let $y=f^{r+k-1}(C_{1})$ be the last point in the cycle. Let $\diamond$ be the preimage of $y$ that is not post-critical. Let $\psi$ be a half twist that swaps $\diamond$ with $f^{r-1}(C_{1})$. Output $f\psi$, the marked set $M\cup\{\diamond\}$ the portrait of $f\psi$ (in this case, not all points in $M$ are vertices in the portrait of $f\psi$), and the mapping class $\phi$. Then $C_{2}$ occurs one position later in the cycle of $C_{1}$ under $f\phi$ than $C_{2}$ occurs in the cycle of $C_{1}$ under $f$, so $q$ increased by one and the cycle length remained constant.
The branched cover $f\phi$ only changes the cycle lengths $q-r$ and $k+r-q$ and no other portrait features of $\Gamma$.
[…]
Figure 2. The function 5(L) increases the position of $C_{2}$ in the orbit of $C_{1}$ and changes no other portrait features.
Proof.
We consider the two cases:
  • (E)
    We first consider the case where $C_{2}$ is not the first point of the cycle of $C_{1}$ (that is $C_{2}\neq f^{r}(C_{1})$). Since $x$ is not a critical value, the set $f^{-1}(x)$ has two points. Therefore only one of the points $f^{-1}(x)$ is post-critical.
    As above, $f^{r-1}(C_{1})$ is the last point in the pre-period of $C_{1}$ under $f$. Likewise, $f^{r-1}(C_{1})$ is the last point of the pre-period of $C_{1}$ under $f\phi$. Moreover, $f\phi$ maps both $f^{r-1}(C_{1})$ and $f^{r}(C_{1})$ to $x$. Then $x$ is the first point of the cycle of $C_{1}$ under $f\phi$. Because $\phi$ is supported on a neighborhood of an arc between $f^{r-1}(C_{1})$ and $\diamond$, the rest of the cycle follows as before. Therefore $C_{2}$ occurs one position earlier in the cycle of $C_{1}$ under $f\phi$ than in the cycle of $C_{1}$ under $f$, and no other portrait features of $\Gamma_{f}$ change under $f\phi$.
  • (L)
    Next we consider the case where $C_{2}$ is not the second last point of the cycle (that is, it does not map to $f^{r+k-1}(C_{1})$). Since $y=f^{r+k-1}(C_{1})$ is not a critical value, the set $f^{-1}(y)$ has two points. Because the $C_{2}$ is in the cycle of $C_{1}$, the portrait for $f$ has at most one pre-period. Therefore only one of the points $f^{-1}(y)$ is post-critical.
    As above, $f^{r-1}(C_{1})$ is the last point in the pre-period of $C_{1}$ under $f$. Likewise, $f^{r-1}(C_{1})$ is the last point of the pre-period of $C_{1}$ under $f\phi$. Moreover, $f\phi$ maps both $f^{r-1}(C_{1})$ and $f^{r+k-2}(C_{1})$ to $y$. Then $y$ is the first point of the cycle of $C_{1}$ under $f\phi$. Because $\phi$ is supported on a neighborhood of an arc between $f^{r-1}(C_{1})$ and $\diamond$, the rest of the cycle follows as before. Therefore $C_{2}$ occurs one position later in the cycle of $C_{1}$ under $f\phi$ than in the cycle of $C_{1}$ under $f$, and no other portrait features of $\Gamma_{f}$ change under $f\phi$.
Function 6 (Move $C_{2}$ from a cycle to the pre-period of $C_{1}$).
Take as input a marked branched cover $f$ with marked set $M$ and associated portrait $\Gamma$ with one component such that one critical point is strictly pre-periodic and the other critical point is periodic. Keeping to our conventions, assume that $C_{2}$ is in the cycle of $C_{1}$, and assume that $C_{2}$ is not the first point in the cycle. As above, let $r$ be the length of the pre-period of $C_{1}$, so that $f^{r-1}(C_{1})$ is the last point of the pre-period. Let $q$ be such that $C_{2}=f^{q}(C_{1})$ (with $q>r$). Let $\phi$ be a half twist that swaps $f^{r-1}(C_{1})$ with $f^{q-1}(C_{1})$. Output the marked branched cover $f\phi$, the marked set $M$, the portrait of $f\phi$, and the mapping class $\phi$. The pre-period of $C_{1}$ under the branched cover $f\phi$ consists of $\{C_{1},f(C_{1}),\cdots,f^{r-1}(C_{1}),C_{2}=f^{q}(C_{1}),\cdots,f^{r+k-1}(C_{1})\}$; see Figure 3. The branched cover $f\phi$ does not change the number of components of $\Gamma_{f}$.
Proof.
The branched cover $f\phi$ maps $f^{r-1}(C_{1})$ to $C_{2}$ and maps $f^{q-1}(C_{1})$ to $f^{r}(C_{1})$. Because $\phi$ can be chosen to be supported on a neighborhood of an arc connecting $f^{q-1}(C_{1})$ and $f^{r-1}(C_{1})$ that is disjoint from the other critical and post-critical points, we have that $(f\phi)^{\ell}(C_{1})=f^{\ell}(C_{1})$ for all $\ell\neq r-1,q-1$.
The pre-period of $C_{1}$ under $f\phi$ is $\{C_{1},f(C_{1}),\cdots,f^{r-1}(C_{1}),C_{2},f^{q+1}(C_{1}),\cdots,f^{r+k-1}(C_{1})\}$. The cycle of $C_{1}$ is $\{f^{r}(C_{1}),\cdots,f^{q-1}(C_{1})\}$. In particular, the portrait of $f\phi$ has one component, but all other portrait features of $f\phi$ may be different than those in the portrait of $f$. ∎
@@
Figure 3. Applying Function 6 creates a new function where $f^{r-1}(C_{1})$ maps to $C_{2}$.
Function 7 (Make a strictly pre-periodic critical point periodic).
Take as input a marked branched cover $f$ with marked set $M$ and associated portrait $\Gamma$ such that (at least) one critical point is strictly pre-periodic. Suppose the critical point $C_{j}$ is strictly pre-periodic under $f$. Let $r$ be the length of the pre-period of $C_{j}$ under $f$, and let $k$ be the length of the cycle of $C_{j}$ under $f$. Note that $C_{j}$ may be in the pre-period of another critical point. Assume that $C_{j}$ is not the image of the other critical point. Then $f^{-1}(C_{j})$ consists of two points, and at most one of them is post-critical. Let $\diamond$ be a preimage of $C_{j}$ that does not lie in the post-critical set of $f$. Let $\phi$ be a half twist that swaps $\diamond$ and $f^{r+k-1}(C_{j})$ (the last point of the cycle of $C_{j}$ under $f$). Output the marked branched cover $f\phi$, the marked set $M\cup\{\diamond\}$, the portrait of $f\phi$, and the mapping class $\phi$. Then $C_{j}$ is in the cycle of $f\phi$ and the cycle has length $k+r$. We refer to the change in the portrait that results from applying $\phi$ as “applying Function 7 to $C_{j}$.” The branched cover $f\phi$ does not change the number of portrait components of $\Gamma_{f}$.
@@
Figure 4. Applying Function 7 to the critical point $C_{j}$ creates a new function in for which $C_{j}$ periodic.
Function 8 (Merge components).
Take as input a marked branched cover $f$ with marked set $M$ and associated portrait $\Gamma$ such the portrait has two (non-empty) components, and further assume that each component has at least two vertices. Call the components $\Gamma_{1}$ and $\Gamma_{2}$ where $\Gamma_{i}$ is the component containing $C_{i}$ for $i\in\{1,2\}$. Let $f^{r_{1}}(C_{1})$ and $f^{r_{2}}(C_{2})$ denote the first element in the cycles of $\Gamma_{1}$ and $\Gamma_{2}$ (here $r_{1}$ or $r_{2}$ can be 0), and let $\phi$ be the half twist that swaps $f^{r_{1}}(C_{1})$ with $f^{r_{2}}(C_{2})$. Output the marked branched cover $f\phi$, the marked set $M$, the portrait of $f\phi$, and the mapping class $\phi$.
[…]
Figure 5. Function 8 may change all portrait features.
The single non-empty component of the portrait of $f\phi$ will have one pre-period for each pre-period of a component of $f$. If $f$ had two pre-periods, then the portrait of $f\phi$ will have two disjoint pre-periods.
Proof of Merge.
The branched cover $f\phi$ maps $f^{r_{1}}(C_{1})$ to $f^{r_{2}+1}(C_{2})$ and maps $f^{r_{2}}(C_{2})$ to $f^{r_{1}+1}(C_{1})$. This implies that $f\phi$ has a cycle of length $k_{1}+k_{2}$, where $k_{1}$ and $k_{2}$ are the cycle lengths of $\Gamma_{1}$ and $\Gamma_{2}$, respectively. Since $\phi$ can be chosen so that it is supported on a neighborhood of an arc between $f^{r_{1}}(C_{1})$ and $f^{r_{2}}(C_{2})$ that is disjoint from all other critical and post-critical points $f\phi$ and $f$ will be equal for all critical and post-critical points besides $f^{r_{1}}(C_{1})$ and $f^{r_{2}}(C_{2})$. Therefore, the critical points in $\Gamma_{1}$ and $\Gamma_{2}$ which are strictly pre-periodic must remain strictly pre-periodic. This implies that the number of pre-periods of $f\phi$ is equal to the number of pre-periods of $\Gamma_{1}$ plus the number of pre-periods of $\Gamma_{2}$, and in the case where $f\phi$ has two pre-periods these pre-periods must be disjoint. ∎
Function 9 (Split portrait from one to two components).
  1. (1)
    1. (a)
      If either $f(C_{1})=C_{2}$ or $f(C_{2})=C_{1}$, let $\phi$ be a half twist that swaps $C_{1}$ with $C_{2}$. Output the marked branched cover $f\phi$, the marked set $M$, the portrait of $f\phi$, and the mapping class $\phi$. The portrait of $f\phi$ has two components, one where one critical point maps to itself, and the other critical point is in a cycle of period $k-1$.
    2. (b)
      Otherwise, let $q$ be such that $f^{q}(C_{1})=C_{2}$. Let $\phi$ be a half twist that swaps $f^{q-1}(C_{1})$ with $f^{k-1}(C_{1})$ (where $k$ is the length of the cycle). Output the marked branched cover $f\phi$, the marked set $M$, the portrait of $f\phi$, and the mapping class $\phi$. The portrait of $f\phi$ has two cyclic components – $C_{1}$ is in a cycle of length $q$, $C_{2}$ is in a cycle of length $k-q$.
  2. (2)
  3. (3)
    Suppose one critical point is in the pre-period of the other. Let $C_{2}$ be in the orbit of $C_{1}$, and let $q$ be such that $f^{q}(C_{1})=C_{2}$. Let $\diamond$ denote a pre-image of $C_{1}$. Let $\phi$ be a half twist that swaps $\diamond$ and $f^{q-1}(C_{1})$ (which could be $C_{1}=f^{0}(C_{1})$). Output $f\phi$, the marked set $M\cup\{\diamond\}$, the portrait of $f\phi$ (in this case not all points of $M$ are vertices of the portrait), and the mapping class $\phi$. Then the portrait of $f\phi$ has two components: $C_{1}$ is in a cyclic component of length $q-1$ and $C_{2}$ is in a component that is the same as the forward orbit of $C_{2}$ under $f$. In particular, $C_{1}$ is cyclic under $f\phi$ and $C_{2}$ is strictly pre-periodic under $f\phi$.
  4. (4)
[…]
    Figure 6. Function 9(4) may change all portrait features of $\Gamma_{f}$
  5. (5)
    Suppose that $C_{1}$ and $C_{2}$ are both strictly pre-periodic and their pre-periods intersect non-trivially. Let $\ell_{1}>0,\ell_{2}\geq 0$ be such that $f^{\ell_{1}}(C_{1})=f^{\ell_{2}}(C_{2})$. Let $\diamond$ denote a pre-image of the first critical point, $C_{1}$. Let $\phi$ be a half twist that swaps $\diamond$ with $f^{\ell_{1}-1}(C_{1})$. Output $f\phi$, the marked set $M\cup\{\diamond\}$, the portrait of $f\phi$ (in this case not all points of $M$ are vertices of the portrait), and the mapping class $\phi$. The portrait $f\phi$ has two components: $C_{1}$ is in a cyclic component of length $\ell_{1}$ and $C_{2}$ is in a component that is the same as the forward orbit of $C_{2}$ under $f$.
Function 9 can change all portrait features of $f$.
[…]
$\ddots$
[…]
$\ddots$
[…]
Figure 7. Function 9(5) may change all portrait features of $\Gamma_{f}$.
@@
(1) In (a), assume up to relabeling that $f(C_{1})=C_{2}$. As defined in the statement of the Function, $\phi$ swaps $C_{1}$ and $C_{2}$. Then $f\phi$ maps $C_{1}$ to itself and maps $C_{2}$ to $f(C_{2})=f^{2}(C_{1})$. Since $\phi$ can be chosen to be supported on a neighborhood of an arc between $C_{1}$ and $C_{2}$, $(f\phi)^{\ell}(C_{1})=f^{\ell}(C_{2})$ for all $12\leq\ell\leq k-$ 1. In particular, the portrait of $f\phi$ has two components: one consisting of the fixed point $C_{1}$, the other is a cyclic component consisting of $\{C_{2},f^{2}(C_{1}),\cdots,f^{k-1}(C_{1})\}$.
(2) The branched cover $f\phi$ maps $f^{r-1}(C_{1})$ to $C_{1}$. Since $\phi$ can be chosen so that it is supported on a neighborhood of an arc containing $\diamond$ and $f^{r-1}(C_{1})$, the maps $f$ and $f\phi$ are equal on all critical and post-critical points except for $f^{r-1}(C_{1})$.
(5) The branched cover $f\phi$ maps $f^{\ell_{1}-1}(C_{1})$ to $C_{1}$. Since $\phi$ can be chosen so that it is supported on a neighborhood of an arc containing $\diamond$ and $f^{\ell_{1}-1}(C_{1})$, $f$ and $f\phi$ are equal on all critical and post-critical points except for $f^{\ell_{1}-1}(C_{1})$. Then $C_{1}$ is periodic under $f\phi$ with period $\{C_{1},f(C_{1}),\cdots,f^{\ell_{1}-1}(C_{1})\}$. The other component of the portrait of $f\phi$ consists of the orbit of $C_{2}$ and the pre-period is the same as the pre-period of $C_{2}$ under $f$ and the cycle is the same as the cycle of $C_{2}$ under $f$. ∎
Function 10 (Create a new pre-period: one (cyclic) component).
Take as input a marked branched cover $f$ with marked set $M$ and associated portrait $\Gamma$ such that the portrait has one (non-empty) component, which is a cycle. That is: both $C_{1}$ and $C_{2}$ are periodic. Let $k$ be the length of the cycle and assume $k\geq 4$. Suppose further that $C_{2}$ is not $f(C_{1})$ and $C_{1}$ is not $f(C_{2})$. Then $C_{2}$ has two preimages under $f$, one is post-critical, the other is not post-critical. Let $\diamond$ be the preimage of $C_{2}$ that is not post-critical. Let $\phi$ be a half twist that swaps $\diamond$ and $f^{k-1}(C_{1})$. Output $f\phi$, the marked set $M$, the portrait of $f\phi$, and the mapping class $\phi$. Then $C_{1}$ is strictly pre-periodic under $f\phi$. The branched cover $f\phi$ does not change the number of portrait components of $\Gamma_{f}$.
[…]
Figure 8. Function 10 adds a pre-period to a cycle containing both critical points.

### v2

@@ 3. Portrait functions
    Suppose that $C_{1}$ and $C_{2}$ are both strictly pre-periodic and their pre-periods are distinct and intersecting. Let $\ell_{1}>0,\ell_{2}\geq 0$ be such that $f^{\ell_{1}}(C_{1})=f^{\ell_{2}}(C_{2})$. Let $\diamond$ denote a pre-image of the first critical point, $C_{1}$. Let $\phi$ be a half twist that swaps $\diamond$ with $f^{\ell_{1}-1}(C_{1})$. Output $f\phi$, the marked set $M\cup\{\diamond\}$, the portrait of $f\phi$ (in this case not all points of $M$ are vertices of the portrait), and the mapping class $\phi$. The portrait $f\phi$ has two components: $C_{1}$ is in a cyclic component of length $\ell_{1}$ and $C_{2}$ is in a component that is the same as the forward orbit of $C_{2}$ under $f$.
Function 1 can change all portrait features of $f$.
@@
Figure 3. Function 1(5) may change all portrait features of $\Gamma_{f}$.
Proof.
(1) In (a), assume up to relabeling that $f(C_{1})=C_{2}$. As defined in the statement of the Function, $\phi$ swaps $C_{1}$ and $C_{2}$. Then $f\phi$ maps $C_{1}$ to itself and maps $C_{2}$ to $f(C_{2})=f^{2}(C_{1})$. Since $\phi$ can be chosen to be supported on a neighborhood of an arc between $C_{1}$ and $C_{2}$, $(f\phi)^{\ell}(C_{1})=f^{\ell}(C_{2})$ for all $2\leq\ell\leq k-1$. In particular, the portrait of $f\phi$ has two components: one consisting of the fixed point $C_{1}$, the other is a cyclic component consisting of $\{C_{2},f^{2}(C_{1}),\cdots,f^{k-1}(C_{1})\}$.
(2) The branched cover $f\phi$ maps $f^{r-1}(C_{1})$ to $C_{1}$. Since $\phi$ can be chosen so that it is supported on a neighborhood of an arc containing $\diamond$ and $f^{r-1}(C_{1})$, the maps $f$ and $f\phi$ agree on all critical and post-critical points except for $f^{r-1}(C_{1})$.
(5) The branched cover $f\phi$ maps $f^{\ell_{1}-1}(C_{1})$ to $C_{1}$. Since $\phi$ can be chosen so that it is supported on a neighborhood of an arc containing $\diamond$ and $^{\ell_{1}-1}(C_{1}), ff$ and $f\phi$ are equal on all critical and post-critical points except for $f^{\ell_{1}-1}(C_{1})$. Then $C_{1}$ is periodic under $f\phi$ with period $\{C_{1},f(C_{1}),\cdots,f^{\ell_{1}-1}(C_{1})\}$. The other component of the portrait of $f\phi$ consists of the orbit of $C_{2}$ and the pre-period is the same as the pre-period of $C_{2}$ under $f$ and the cycle is the same as the cycle of $C_{2}$ under $f$. ∎
Function 2 (Make a strictly pre-periodic critical point periodic).
Take as input a marked branched cover $f$ with marked set $M$ and associated portrait $\Gamma$ with two non-empty components such that (at least) one critical point is strictly pre-periodic. Suppose the critical point $C_{j}$ is strictly pre-periodic under $f$. Let $r$ be the length of the pre-period of $C_{j}$ under $f$, and let $k$ be the length of the cycle of $C_{j}$ under $f$. Because $C_{1}$ and $C_{2}$ are in different components, neither critical point is the image of the other. Therefore $f^{-1}(C_{j})$ consists of two points, and at most one of them is post-critical. Let $\diamond$ be a preimage of $C_{j}$ that does not lie in the post-critical set of $f$. Let $\phi$ be a half twist that swaps $\diamond$ and $f^{r+k-1}(C_{j})$ (the last point of the cycle of $C_{j}$ under $f$). Output the marked branched cover $f\phi$, the marked set $M\cup\{\diamond\}$, the portrait of $f\phi$, and the mapping class $\phi$. Then $C_{j}$ is in the cycle of $f\phi$ and the cycle has length $k+r$. We refer to the change in the portrait that results from applying $\phi$ as “applying Function 2 to $C_{j}$.” The branched cover $f\phi$ does not change the number of portrait components of $\Gamma_{f}$.
@@
Figure 4. Applying Function 2 to the critical point $C_{j}$ creates a new function in for which $C_{j}$ periodic.
Function 3 (Decrease a cycle length).
Take as input a marked branched cover $f$ with marked set $M$ and portrait $\Gamma$. We will assume that the portrait of $f$ consists of two cyclic components. If the length of the cycle of $C_{j}$ is greater than one, let $f^{\ell}(C_{j})$ and $f^{\ell+1}(C_{j})$ be (distinct) post-critical points in the cycle of $C_{j}$. The points $f^{\ell}(C_{j})$ and $f^{\ell+1}(C_{j})$ will not both be critical points because each critical point is in a different component. Then let $\phi$ be a half twist that swaps $f^{\ell}(C_{j})$ and $f^{\ell+1}(C_{j})$. Output the marked branched cover $f\phi$, the marked set $M$, the portrait of $f\phi$, and the mapping class $\phi$. The cycle of $C_{j}$ in the portrait of $f\phi$ has length one less than the the length of the cycle of $C_{j}$ in the portrait of $f$.
The branched cover $f\phi$ only affects the length of the cycle, and no other portrait features of $\Gamma_{f}$. We refer to the change in the portrait that results from applying $\phi$ as “applying Function 3 to the point $f^{\ell}(C_{j})$.”
@@
Proposition 3.1.
The set of abstract portraits that are compatible with a quadratic branched cover are closed under Functions 1-3.